Two industrial cleaning products salesman A and B, must each make two calls per day, one in the morning and one in the afternoon. Salesman A has a probability of 0.45 of selling a cleaner on any call, while salesman B (a novice) has a probability of 0.2 of a sale. The two salesman work independently of each other, and each salesman’s morning and afternoon results are independent of each other.
3.*Question* :
Find the probability that salesman B makes no sales (rounded off to two decimals)?

Answer by ikleyn(52864)

.

P = (1-0.2)*(1-0.2) = 0.8*0.8 = 0.64.


The formula is self-explanatory.